# SQL-Factory: A Multi-Agent Framework for High-Quality and Large-Scale SQL Generation

### v4

@@ 5.1.3. Baseline
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both random generation methods and learning-based generation methods. Specifically, we include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs over a given schema, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale queries. Additionally, we exclude template-based methods (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018) from our comparison, as they rely heavily on predefined templates and are not suitable for generating massive and structurally diverse SQL workloads.
Text-to-SQL Baselines. For evaluating the effectiveness of the generated data on downstream Text-to-SQL tasks, we fine-tune two base models: Qwen2.5-Coder-1.5B and Qwen2.5-Coder-3B, using SQL-NL pairs derived from the SQL queries synthesized by SQL-Factory. The NL annotations are obtained via a post-processing step, as described in Section 5.3. We compare the performance of these fine-tuned models with several model of similar scale, including Codes-1B, and Codes-3B (Li et al., 2024). This setup allows us to examine whether augmenting training with generated queries generated by SQL-Factory can improve model performance.
Query Clustering Baselines. For the SQL clustering task, we evaluate several established similarity models. These include statistical and heuristic approaches such as Aouiche (Aouiche et al., 2006), Aligon (Aligon et al., 2014), and Makiyama (Makiyama et al., 2015), the pretrained model-based PreQR, and two general-purpose language models: BERT-Base and BERT-Large (Devlin et al., 2019). We further include versions of BERT fine-tuned with contrastive learning on SQL-Factory data to highlight the effectiveness of our generated queries in improving SQL representation learning.
@@ 5.1.4. Implementation Details
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source model optimized for code and SQL-related instruction following capability. It is served using vLLM to support efficient batch decoding at scale, enabling rapid query expansion. It is also use a temperature of 0.8 for diversity.
@@ 5.2.1. Overview of Generated Queries
Key aspects such as join depth, predicate usage, subquery nesting, aggregate usage, and overall token length provide meaningful insights into the expressiveness and structural diversity of SQL queries. These dimensions serve as proxies for measuring the coverage of different logical constructs and the capability to reflect real-world query distributions. Therefore, we conduct a comprehensive analysis of the SQL queries along these dimensions in Figure 6.
Figure 6(a) illustrates the distribution of join operations per query. Notably, 78.6% of all queries involve at least one join, and 30% contain two or more join. This indicates that SQL-Factory effectively generates multi-table queries, showcasing its capacity for relational reasoning and schema-aware generation. Figure 6(b) analyzes the number of predicates in queries. More than 58% of the generated queries contain two or more predicates, with 9.8% having four or more. This confirms that SQL-Factory not only generates structurally rich queries but also embeds fine-grained filtering logic, which is critical for supporting meaningful data retrieval and realistic workload simulation. Besides, Figure 6(c) and Figure 6(d) present the proportions of queries that contain nested subqueries and aggregation operations, respectively. These results demonstrate that a large portion of the generated SQLs utilize advanced SQL constructs, further validating the model’s ability to produce structurally rich and expressive queries. Finally, Figure 6(e) reports the distribution of token lengths across these queries. The x-axis represents the number of tokens for each query, while the y-axis shows the corresponding frequency. The distribution reveals a broad range of query lengths, confirming that SQL-Factory does not produce templated or overly uniform queries, but rather supports diverse query formulations at varying levels of complexity.
Overall, these results confirm that SQL-Factory produces syntactically diverse and structurally expressive SQL queries, with rich combinations of logical constructs. Such diversity is essential for training robust models in downstream applications.
@@ 5.2.2. Cost Analysis of SQL Generation
Benchmark # DB # Table # Col/DB # Queries Price (\$)
TPC-DS (Nambiar and Poess, 2006) 1 24 449 50,000 51.78
IMDB (Maas et al., 2011) 1 21 108 40,000 18.47
Spider (Yu et al., 2018) 69 795 26.82 130,000 42.16
BIRD (Li et al., 2023a) 146 524 54.56 160,000 67.50
Table 1 summarizes the key statistics of SQL queries generated by SQL-Factory across four benchmarks. These include the number of databases, tables, average number of columns per database, the total number of generated queries, and the corresponding cost of API.
The four benchmarks span a diverse range of schema structures. TPC-DS and IMDB each consist of a single database but vary substantially in complexity, TPC-DS features 449 columns while IMDB has 108 specifically. Meanwhile, BIRD and Spider contain hundreds of databases with fewer columns per schema on average, representing more fragmented but wider coverage scenarios. Across these benchmarks, we observe that the total number of generated queries generally correlates with the schema’s complexity and scale. For example, the largest numbers of queries are generated for BIRD (160,000) and Spider (130,000), which contain the most tables and databases. In contrast, IMDB yields 40,000 SQLs due to its simpler, single-schema structure. This adaptive generation behavior emerges naturally from our multi-agent framework, which allocates more generation rounds and expansions in response to richer or more varied schema coverage requirements.
Despite the varying generation needs, SQL-Factory maintains high efficiency. The total cost across all four benchmarks remains under \$200. Notably, if the entire workload were generated solely by GPT-4o without the Expansion Team, the API cost would approach \$1000, highlighting the substantial cost savings enabled by our hybrid design. Together, these results demonstrate that SQL-Factory is capable of producing SQL workloads at scale, while maintaining both schema adaptivity and cost-effectiveness.
@@ 5.2.3. Comparative Analysis on Query Diversity
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: SQLSmith (Seltenreich, 2016), a grammar-driven random query generator that walks through SQL syntax trees; LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate queries through a finite-state abstraction; and OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of SQL queries.
@@
Hybrid Similarity. We use our proposed hybrid similarity metric as described in Section 4.1.2 to measure the average pairwise similarity between queries. This metric integrates token-level, structural-level, and embedding-level similarity signals, providing a robust measure of redundancy in query content and structure. Lower scores indicate higher diversity. As shown in Table 2, SQL-Factory achieves a balanced similarity score of 0.637, significantly lower than LearnedSQLGen (0.863) and OmniSQL (0.688), suggesting that it avoids over-repetition while maintaining meaningful structure. Besides, the low similarity score for SQLSmith stems largely from its use of excessive syntactic randomness. It frequently inserts arbitrary user-defined functions, opaque predicates, and irrelevant clauses that contribute to lexical variance but not to meaningful semantic or structural diversity.
Vendi Score. The Vendi Score (Friedman and Dieng, 2023) is a recently proposed metric for evaluating diversity based on the exponential of the entropy of the eigenvalues of a similarity matrix computed from the data. It does not rely on predefined labels or reference distributions, making it especially suitable for open-ended generation tasks. We utilize Vendi Score to assess the diversity of our generated datasets with two embedding backbones: SimCSE (Gao et al., 2021), a general-purpose sentence encoder trained via contrastive learning, and SQL-Encoder (Pourreza et al., 2024), a specialized model pretrained for SQL embeddings. Higher Vendi scores indicate greater overall diversity in the embedding space. As shown in Table 2, SQL-Factory achieves the highest Vendi Score under both encoders. In contrast, SQLSmith demonstrates low similarity but also fails to produce coherent queries. LearnedSQLGen suffers from high redundancy due to its pattern imitation and optimization-driven training. Besides, OmniSQL is limited by its monolithic generation strategy.
@@ 5.2.4. Schema-Aware Query Allocation Analysis
To evaluate whether SQL-Factory enables schema-aware query planning, we compare it with OmniSQL (Li et al., 2025a) across two key metrics: the number of queries for each table and the structural complexity of each table. We define a table’s complexity score as the sum of its number of columns and twice the number of foreign keys, which reflects both structural richness and relational connectivity. Figure 7 and Figure 8 illustrate these distributions for OmniSQL and SQL-Factory, respectively.
Figure 7(a) shows that exhibits a highly skewed distribution, where a small number of tables dominate the workload while most tables receive very few queries. However, as shown in Figure Figure 7(b), the distribution of table complexity in the same benchmark is much more balanced, with the majority of tables falling in a moderate complexity range. This misalignment suggests that OmniSQL does not account for schema structure when allocating queries, which may result in under-representation of complex tables.
In contrast, SQL-Factory produces a query distribution that more closely matches schema complexity. Figure 8(a) shows that most tables receive a moderate number of queries, and this pattern aligns well with the table complexity distribution in Figure 8(b), where moderately complex tables are also the most common. This consistency indicates that SQL-Factory can allocate queries in a way that reflects the underlying schema properties, supporting schema-aware planning and improving the quality of the workload.
@@ 5.2.5. Generation Stopping Criteria
Although SQL-Factory is capable of generating an unbounded number of SQL queries for any benchmark, it is also important to ensure high overall data quality. To this end, we design a generation termination policy that avoids excessive accumulation of redundant or low-diversity queries.
To determine when to terminate for each benchmark, we monitor two key metrics: (1) the expansion ratio, defined as the proportion of all queries in the SQL Pool that originate from the Expansion Team, and (2) the average hybrid similarity, which reflects query redundancy in the overall pool.
As illustrated in Figure 9, the system is initialized by the Generation Team, which produces a small number of diverse seed queries. After this warm-up phase, the Expansion Team begins to take over, resulting in a rise in both the expansion ratio and similarity across the SQL Pool. This indicates that a growing number of queries are being synthesized based on previously explored patterns.
To maintain overall query diversity, the Management Agent periodically switches back to the Generation Team. These transitions manifest as synchronized dips in both expansion ratio and similarity on the curve, reflecting the injection of more novel queries. This alternating pattern of generation and expansion continues iteratively, with diminishing gains over time. Once the Expansion Team becomes dominant and the similarity curve shows a sustained upward trend, it indicates that the SQL Pool has reached semantic saturation, prompting us to terminate the synthesis process.
For example, after synthesizing approximately 50,000 queries for TPC-DS, we observe that the Expansion Team begins to dominate the process. Specifically, the expansion ratio rapidly decrease while the average similarity among queries steadily rising. This trend indicates that the SQL Pool has become saturated. Continuing the process beyond this point leads to diminishing returns in diversity and semantic coverage. Therefore, we empirically adopt this saturation point as the stopping criterion for each benchmark.
@@ 5.3. Text-to-SQL Task
Table 3. Results on Text-to-SQL Task.
@@
Executability Accuracy Executability Accuracy
Avg Avg Simple Moderate Challenging Avg Simple Moderate Challenging Avg
Codes-1B 90.43 68.09 89.30 81.03 83.45 86.25 42.38 26.51 22.07 35.66
Codes-3B 94.97 75.05 92.32 85.34 88.28 89.83 53.30 33.19 28.97 44.92
Qwen2.5-Coder-1.5B 85.69 69.15 63.24 49.78 46.21 57.56 34.92 19.83 15.86 28.55
Qwen2.5-Coder-3B 92.84 76.21 73.51 63.15 56.55 68.77 43.78 29.74 17.93 37.09
Qwen2.5-Coder-1.5B+SQL-Factory 97.20 76.50 87.89 74.57 66.21 81.81 50.92 26.94 19.31 40.68
Qwen2.5-Coder-3B+SQL-Factory 97.29 79.69 92.97 82.11 75.86 88.07 55.35 33.62 26.90 46.09

### v5

@@ 5.1.3. Baseline
SQL Generation Baselines. To compare the quality and diversity of generated SQL queries, we consider both random generation methods and learning-based generation methods. Specifically, we include SQLSmith (Seltenreich, 2016), a grammar-driven query generator that produces random SQLs, and two representative learning-based methods: LearnedSQLGen (Zhang et al., 2022), which learns to generate queries via reinforcement learning, and OmniSQL (Li et al., 2025a), which leverages large language models to synthesize large-scale queries. Additionally, we exclude template-based methods (Nambiar and Poess, 2006; Poess and Floyd, 2000; Leis et al., 2018) from our comparison, as they rely heavily on predefined templates and are not suitable for generating massive and structurally diverse SQL workloads.
Text-to-SQL Baselines. To assess the effectiveness of the generated training data, we evaluate a series of commonly used models, including Qwen2.5-Coder-1.5B/3B (Hui et al., 2024), LLaMA3.2-1B/3B (AI, 2024), and Deepseek-Coder-1.3B (Guo et al., 2024), under three training settings: (1) the original model, (2) fine-tuning using the official benchmark training data (e.g., Spider or BIRD), and (3) fine-tuning using the data produced by our SQL generation pipeline.
Query Clustering Baselines. For the SQL clustering task, we evaluate several established similarity models. These include statistical and heuristic approaches such as Aouiche (Aouiche et al., 2006), Aligon (Aligon et al., 2014), and Makiyama (Makiyama et al., 2015), the pretrained model-based PreQR, and two general-purpose language models: BERT-Base and BERT-Large (Devlin et al., 2019).
@@ 5.1.4. Implementation Details
Expansion Agent. For the Expansion Agent, we utilize Qwen2.5-Coder-14B (Hui et al., 2024), a locally deployed open-source model optimized for code and SQL-related instruction following capability. It is served using vLLM to enable rapid query expansion. It is also use a temperature of 0.8 for diversity.
@@ 5.2.1. Overview of Generated Queries
We analyze the generated SQL queries across several dimensions, including join depth, predicate usage, subquery nesting, aggregate usage, and token length. These metrics serve as proxies for expressiveness, structural diversity, and alignment with real-world query patterns.
Figure 6(a) illustrates the distribution of join operations per query. Notably, 78.6% of all queries involve at least one join, and 30% contain two or more join. This indicates that SQL-Factory effectively generates multi-table queries, showcasing its capacity for relational reasoning and schema-aware generation. Figure 6(b) analyzes the number of predicates in queries. More than 58% of queries contain two or more predicates, with 9.8% having four or more. This confirms that SQL-Factory not only generates structurally rich queries but also embeds fine-grained filtering logic, which is critical for supporting meaningful data retrieval and realistic workload simulation. Besides, Figure 6(c) and Figure 6(d) reveal substantial use of nested subqueries and aggregation, confirming the system’s capability to produce advanced constructs. Furthermore, Figure 6(e) reports the distribution of token lengths. The x-axis represents the number of tokens for each query, while the y-axis shows the corresponding frequency. The distribution reveals a broad range of query lengths, confirming that SQL-Factory does not produce templated or overly uniform queries, but rather supports diverse query formulations at varying levels of complexity. In addition, we analyze the proportion of queries that return empty result sets. Only 15.34% of the queries fall into this category, suggesting that most generated queries are not only syntactically valid but also semantically meaningful.
Overall, these results confirm that SQL-Factory produces syntactically diverse and structurally expressive SQL queries. Such diversity is essential for training robust models in downstream applications.
@@ 5.2.2. Cost Analysis of SQL Generation
Benchmark # DB # Table # Col/DB # Queries Price (\$) Time (h)
TPC-DS 1 24 449 50,000 51.78 56.5
IMDB 1 21 108 40,000 18.47 48.2
SPIDER 69 795 26.82 130,000 42.16 233.3
BIRD 146 524 54.56 160,000 67.50 283.1
Table 1 summarizes the key statistics of SQL queries generated by SQL-Factory across four benchmarks. These include the number of databases, tables, average number of columns per database, the total number of generated queries, and the corresponding cost of API. Across these benchmarks, we observe that the total number of generated queries generally correlates with the schema’s complexity and scale. For example, the largest numbers of queries are generated for BIRD (160,000) and Spider (130,000), which contain the most tables and databases. In contrast, IMDB yields 40,000 SQLs due to its simpler, single-schema structure. This adaptive generation behavior emerges naturally from our multi-agent framework, which allocates more generation rounds and expansions in response to richer or more varied schema coverage requirements.
Despite the varying generation needs, SQL-Factory maintains high efficiency. The total cost across all four benchmarks remains under \$200. Notably, if the entire workload were generated solely by GPT-4o without the Expansion Team, the API cost would approach \$1000, highlighting the substantial cost savings enabled by our hybrid design. Besides, generating the first 10,000 queries typically takes around 7 to 11 hours, depending on schema complexity and prompt length. As the number of existing queries grows, operations such as statistical analysis and hybrid similarity computation introduce additional overhead, leading to a gradual slowdown. Nevertheless, the overall throughput remains significantly higher than manual query construction. Together, these results demonstrate that SQL-Factory is capable of producing SQL workloads at scale, while maintaining both schema adaptivity and cost-effectiveness.
@@ 5.2.3. Comparative Analysis on Query Diversity
We conduct a comparative study to assess the query diversity achieved by SQL-Factory against three representative SQL generation methods: (1) SQLSmith (Seltenreich, 2016), a grammar-driven random query generator that walks through SQL syntax trees; (2) LearnedSQLGen (Zhang et al., 2022), which employs a reinforcement learning framework guided by cost-based feedback to generate queries through a finite-state abstraction; and (3) OmniSQL (Li et al., 2025a), a LLM-based framework that leverages prompt engineering to generate millions of SQL queries.
@@
Hybrid Similarity. We use our proposed hybrid similarity metric as described in Section 4.1.2 to measure the average pairwise similarity between queries. As shown in Table 2, SQL-Factory achieves a balanced similarity score of 0.637, significantly lower than LearnedSQLGen (0.863) and OmniSQL (0.688), suggesting that it avoids over-repetition while maintaining meaningful structure. Besides, the low similarity score for SQLSmith is due to its excessive syntactic randomness. It was originally designed for database stress testing, focusing on exposing engine-level bugs rather than generating semantically coherent queries. While its queries increase lexical variance, they often contain irregular constructs and unrealistic patterns that do not align with practical workloads.
Vendi Score. The Vendi Score (Friedman and Dieng, 2023) is a recently proposed metric for evaluating diversity based on the exponential of the entropy of the eigenvalues of a similarity matrix computed from the data. It does not rely on predefined labels or reference distributions, making it especially suitable for open-ended generation tasks. We utilize Vendi Score to assess the diversity of our generated datasets with two embedding backbones: SimCSE (Gao et al., 2021), a general-purpose sentence encoder trained via contrastive learning, and SQL-Encoder (Pourreza et al., 2024), a specialized model pretrained for SQL embeddings. Higher Vendi scores indicate greater overall diversity. As shown in Table 2, SQL-Factory achieves the highest Vendi Score under both encoders. In contrast, SQLSmith demonstrates low similarity but also fails to produce coherent queries. LearnedSQLGen suffers from high redundancy due to its pattern imitation and optimization-driven training. Besides, OmniSQL is limited by its monolithic generation strategy.
5.2.4. Ablation Study on Agent Decomposition
To evaluate the benefits of agent decomposition, we compare SQL-Factory with two single-agent baselines on the TPC-DS benchmark: (1) direct generation using GPT-4o, and (2) direct generation using Qwen2.5-Coder-14B. Each method generates 10,000 queries. As shown in Table 3, SQL-Factory achieves the lowest hybrid similarity and highest Vendi Scores, indicating better structural and semantic diversity. Unlike single-agent approaches, SQL-Factory combines seed generation with targeted expansion, avoiding mode collapse and improving output quality. In addition, SQL-Factory is more cost-efficient, with a significantly lower API cost and moderate time consumption, demonstrating better scalability for large-scale query generation.
Table 3. Comparison with single-agent baselines.
Method Time Price Hybrid Vendi Score Vendi Score
Similarity SQLEncoder SimCSE
GPT-4o 9.2h 104.6\$ 0.672 3.84 4.44
Qwen-14B 23.7h - 0.733 2.70 3.88
SQL-Factory 11.4h 15.22\$ 0.637 4.10 7.94
@@ 5.2.5. Schema-Aware Query Allocation Analysis
To evaluate SQL-Factory’s ability to support schema-aware query planning, we compare it with OmniSQL (Li et al., 2025a) using two metrics: the number of queries per table and each table’s structural complexity. We define table complexity following established schema complexity metrics in database engineering literature (Piattini et al., 2001; Kotsyuba et al., 2021; Pavlic et al., 2008). Specifically, we incorporate three widely recognized structural factors: the number of attributes (NA), the referential degree (RD), and the depth of the referential tree (DRT). The final complexity is calculated as:
(12) $\text{Complexity}(T)=\text{NA}(T)+2\ \text{RD}(T)+\text{DRT}(T).$
This formulation captures both local structural richness and broader relational connectivity. Figure 7 and Figure 8 illustrate the resulting complexity distributions for OmniSQL and SQL-Factory, respectively. Figure 7(a) shows that exhibits a highly skewed distribution, where a small number of tables dominate the workload while most tables receive very few queries. However, as shown in Figure 7(b), the distribution of table complexity in the same benchmark is much more balanced, with the majority of tables falling in a moderate complexity range. This misalignment suggests that OmniSQL does not account for schema structure when allocating queries, which may result in under-representation of complex tables. In contrast, SQL-Factory produces a query distribution that more closely matches schema complexity. Figure 8(a) shows that most tables receive a moderate number of queries, and this pattern aligns well with the table complexity distribution in Figure 8(b), where moderately complex tables are also the most common. This consistency indicates that SQL-Factory can allocate queries in a way that reflects the underlying schema properties, supporting schema-aware planning and improving the quality of the workload.
@@ 5.2.6. Generation Stopping Criteria
Although SQL-Factory can generate an unlimited number of SQL queries for any benchmark, it is also important to ensure high overall data quality. To this end, we employ a termination policy that avoids excessive accumulation of redundant or low-diversity queries. Specifically, we monitor two key metrics: (1) the expansion ratio, defined as the proportion of queries in the SQL Pool that originate from the Expansion Team, and (2) the average hybrid similarity, which reflects overall similarity.
As illustrated in Figure 9, the system is initialized by the Generation Team, which produces a small number of diverse seed queries. After this warm-up phase, the Expansion Team begins to take over, resulting in a rise in both the expansion ratio and similarity across the SQL Pool. This indicates that a growing number of queries are being synthesized based on previously explored patterns. To maintain overall query diversity, the Management Agent periodically switches back to the Generation Team. These transitions manifest as synchronized dips in both expansion ratio and similarity on the curve, reflecting the injection of more novel queries. This alternating pattern of generation and expansion continues iteratively, with diminishing gains over time. Once the Expansion Team becomes dominant and the similarity curve shows a sustained upward trend, it indicates that the SQL Pool has reached semantic saturation, prompting us to terminate the synthesis process.
For example, after synthesizing approximately 50,000 queries for TPC-DS, the Expansion Team begins to dominate the process. Specifically, the expansion ratio rapidly decrease while the average similarity among queries steadily rising. Continuing the process beyond this point leads to diminishing returns in diversity and semantic coverage. Therefore, we empirically adopt this saturation point as the stopping criterion for each benchmark.
@@ 5.3. Text-to-SQL Task
Table 4. Results on Text-to-SQL Task (%). ”+Official-Data” refers to fine-tuning using the original training sets provided by the benchmarks. ”+SQL-Factory” refers to fine-tuning using training data generated by our SQL-Factory framework.
@@
Executable Accuracy Executable Accuracy
Qwen-1.5B 85.7 69.2 57.6 28.6
    +Official-Data 91.9 ($\uparrow$6.2) 73.2 ($\uparrow$4.0) 79.1 ($\uparrow$21.5) 39.2 ($\uparrow$10.6)
    +SQL-Factory 97.2 ($\uparrow$11.5) 76.5 ($\uparrow$7.3) 82.1 ($\uparrow$24.5) 40.9 ($\uparrow$12.3)
Qwen-3B 92.8 76.2 68.8 37.1
    +Official-Data 94.8 ($\uparrow$2.0) 77.8 ($\uparrow$1.6) 88.7 ($\uparrow$19.9) 45.7 ($\uparrow$8.6)
    +SQL-Factory 97.3 ($\uparrow$4.5) 79.7 ($\uparrow$3.5) 88.1 ($\uparrow$19.3) 46.1 ($\uparrow$9.0)
LLama-1B 73.6 41.3 43.3 11.3
    +Official-Data 80.1 ($\uparrow$6.5) 60.3 ($\uparrow$19.0) 62.5 ($\uparrow$19.2) 26.0 ($\uparrow$14.7)
    +SQL-Factory 94.3 ($\uparrow$20.7) 70.6 ($\uparrow$29.3) 71.4 ($\uparrow$28.1) 31.4 ($\uparrow$20.1)
LLama-3B 89.0 60.4 73.9 30.1
    +Official-Data 91.8 ($\uparrow$2.8) 72.2 (11.8) 82.0 ($\uparrow$8.1) 38.9 ($\uparrow$8.8)
    +SQL-Factory 98.7 ($\uparrow$9.7) 76.8 (16.4) 84.7 ($\uparrow$10.8) 42.1 ($\uparrow$12.0)
Deepseek-1.3B 87.2 50.2 63.3 20.3
    +Official-Data 88.7 ($\uparrow$1.5) 70.7 ($\uparrow$20.5) 80.3 ($\uparrow$17.0) 36.9 ($\uparrow$16.6)
    +SQL-Factory 98.3 ($\uparrow$11.1) 75.7 ($\uparrow$25.5) 79.8 ($\uparrow$16.5) 38.3 ($\uparrow$18.0)